# Theory of Mind Enhances Collective Intelligence

### v1

@@ Abstract
Collective Intelligence plays a central role in a large variety of fields, from economics and evolutionary theory to neural networks and eusocial insects, and it is also core to much of the work on emergence and self-organisation in complex systems theory. However, in human collective intelligence there is still much more to be understood in the relationship between specific psychological processes at the individual level and the emergence of self-organised structures at the social level. Previously psychological factors have played a relatively minor role in the study of collective intelligence as the principles are often quite general and applicable to humans just as readily as insects or other agents without sophisticated psychologies. In this article we emphasise, with examples from other complex adaptive systems, the broad applicability of collective intelligence principles while the mechanisms and time-scales differ significantly between examples. We contend that flexible collective intelligence in human social settings is improved by our use of a specific cognitive tool: our Theory of Mind. We identify several key characteristics of psychologically mediated collective intelligence and show that the development of a Theory of Mind is a crucial factor distinguishing social collective intelligence from general collective intelligence. We then place these capabilities in the context of the next steps in artificial intelligence embedded in a future that includes an effective human-AI hybrid social ecology.
@@ 1 Introduction
Collectives are capable of achieving things that individuals alone cannot. Notwithstanding the simplicity or complexity of the individuals, their aggregate behaviour can often be understood as a complex processing of information that individuals store, modify, and transfer between each other producing ‘useful’ collective behaviour at the scale of the whole collective. In most instances of Collective Intelligence (CI), where the agents might be ants in an ant colony, bees in a beehive, or neurons in a neural network, the individual is not aware of the drivers of their behaviour or the behaviour of other agents. For example, a single neuron is neither aware of its own internal processes nor that of a neuron it is connected to, nor is it aware of the end goal to which its activity contributes. Despite both this lack of awareness and the lack of a centralised controller, evolutionary and learning processes have produced an intricate, precise, and highly adaptive system that is capable of functional behaviour that would be impossible for any single neuron to achieve. In other instances of CI, such as teams of humans, or businesses interacting in economic markets, the agents themselves may be highly complex and exhibit varying degrees of purposefulness and awareness. Within this context, we draw attention to the role of psychological factors in improving the CI of human social collectives and quantifying the intelligence of social collectives, both natural and artificial.
@@
We can also compare the competencies of neural cells in networks to the individual competencies of ants in an ant colony. In a recent study [56] it was shown that social structures of some ant colonies are conserved between species that are separated by more than 100 million years of evolution. In the five species studied by Kay et al. [56], they found two social clusters and similarities in the division of labour that are preserved between the species. In a different study, Richardson et al. [91] showed that individuals within an ant colony play an important leadership role and that the behaviour of these individuals significantly improved the collective performance of the ants. Ants are also capable of changing their social structure in the event of pathogenic infestation of their colony. In a 2021 article, Stockmaier and colleagues [102] review the research on social distancing and other social restructuring that occurs with conspecifics in order to reduce the impact of pathogens by changing their social cues, signals, and other behaviours for the collective benefit of the colony. These two very different systems, neural networks and ant colonies, are examples of complex collective intelligences where the individuals (neurons, ants) are complex in their own right, but they signal each other in order to restructure their relationships so as to adapt their collective competencies to external signals. The neural networks are prototypical solid brains and ant colonies are prototypical liquid brains.
@@
In this article we use information theory to quantify how much computation in a CI is ‘emergent’ and how much is simply independent information processing by single agents. In general, we wish to capture the notion of the whole (computational process) being greater than the sum of the (independent) parts. We translate this to the simple notion that to the extent to which this inequality holds: Whole - ($\sum$(∑ (Parts)>0)>0) > 0 is the extent to which we will say a system exhibits non-trivial CI, noting that there are multiple possible implementations of this approach [54]. The Parts is how much computation a single agent is carrying out from one time step to the next such that the sum is the total of all agents’ independent computations. The Whole is the totality of computation in the system, it includes all single agent computations, pairwise computations, and higher order interactions between agents. Our measure will not be unique in any of its specifics, but it serves to quantify the CI of a system for comparative analysis. This approach also has much in common with that of Moore et al. [84] in which information theory is used to measure the collective intelligence in biological systems.
@@
In this work we will ask an analogous question of human agents: What is there in human psychology that allows us to learn to read, interpret, and re-write our interpersonal information content in a way that allows predictive control over our collective behaviour? We will not be able to explore all of the possible interpretations of this question here, but we posit that our Theory of Mind (ToM) is a suite of cognitive skills that allows individuals to have goal directed control over collective outcomes. Originally ToM was used to describe our ability to infer the unobserved mental states of other people [34] such as desires and beliefs, an ability humans are particularly good at and other animals much less so [86, 59]. But recently it has been shown that ToM is predictive of group performance as well [121, 31], empirically demonstrating the role of ToM in going beyond representations of the internal states of others to using that knowledge in a social setting to improve the collective outcomes for the group. In order to model ToM in a tractable fashion, we will focus on the narrower game theory of mind [123], and the Beliefs, Preferences, and Constraints (BPC) interpretation of game-theoretic decisions put forward by Gintis [35]. In this approach, what agents understand of other agents’ hidden states are the BPC that structure their observable behaviours.
@@
We will consider this question in the framework of agent interactions that extend agent utilities in a simple but novel way. We quantify our results using information theory to show the impact that a correctly deployed ToM has to direct agents’ behaviours in order to increase our CI. The models are simple but they illustrate the central notion that understanding the “beliefs, preferences, and constraints” [36, 37] of others can be used to improve the CI of a complex social system.
In Section 2, we describe the liquid–solid dichotomy of interacting agents, review extant models of ToM, and provide perspective on the interplay between social network structures and ToM. In Section 3, we provide illustrative examples supporting different aspects of our argument, introducing our measure of computation and applying it to a simple empirical example. In Section 4 we review the psychology of social fluidity and the variety of social outcomes that this fluidity makes possible. We also use a simple multi-agent system to describe how a ToM can be used to improve the computational processes, i.e. the CI, of interacting agents. Finally, in Section 5, we discuss the broader implications of this approach.
@@ 2.1 Network topology and the “Solid Brain, Liquid Brain” framework
Whole classes of models that describe which biological or artificial structures are capable of some form of computation, either in potentia or in practice, can often be usefully represented using morphospaces [16]. Computational morphospaces [101] have proven effective in studying key properties of complex adaptive systems, for example the statistical mechanics of information processing and structural variations [2, 4]. This sheds light on how energy constraints influence the evolution and adaptability of neural networks across a variety of biological systems. Arsiwalla et al. [2] have examined how liquid brains and solid brains fit within such a framework, by comparing the flexibility and adaptability of different neural architectures [2]. The dimensions of their framework are three different types of complexity that a system may display: autonomic, computational, and social complexity. We suggest that other axes for consideration are a system’s solid–liquid dimension as Solé et al. [101] and Ollé-Vila et al. [85] have done, as well as the system’s degree of ToM (see Section 2.2) and the systems information processing capacity (see Section 3).
@@ 2.2 Models of another agent’s internal states
In the field of artificial intelligence, Jara-Ettinger [53] proposed the use of Inverse Reinforcement Learning (IRL) as a model of agential ToM, whereby agents modelling other agents’ mental states is equivalent to inferring an unobserved world model the other agent uses in their decision-making, as well as their reward function. Jara-Ettinger discusses some key limitations of IRL, such as the difficulty in recovering an agent’s beliefs and desires even while assuming that all agents are identical in their choice-making. IRL has been implemented in many different algorithmic forms, and their applicability as a basis for ToM was recently reviewed by Ruiz-Serra and Harré [93].
@@
This leads to another core finding related to ToM: the interplay between language and ToM in humans. Chomsky has noted both how well developed this ability is in humans and how communication sits in relation to our mental states [18, p. 10]:
@@
Language use is related to this causal reasoning about hidden variables via the representational view of language due to their intersection with ToM [32]. In the representational view of language people use specific grammatical structures to represent complex events and then to reason from them [23]. These structures serve as a cognitive tool, particularly in representing others’ mental states, enabling the expression of false beliefs, lies, or mistakes. It has also been shown that these are strong predictors of children’s false belief understanding, a canonical test of ToM, and targeted syntactical training can improve false belief reasoning [24]. Notably, in developmental learning, language has a stronger influence on ToM abilities than vice versa [80]. From this we identify causal reasoning about hidden cognitive states with causal reasoning about the behaviour of others via a shared representation of our social environment, mediated by specific syntactical structures. We discuss shared representations and their use further in the next section.
@@ 3.2 A triadic example of evolutionary learning at long time scales
This type of second order interaction occurs in other ecological examples as well [38] where Trait-Mediated Indirect Interactions (TMIIs) induce hyper-graphs of interactions between species. More complicated interactions have also been observed in which a plant that is being attacked emits infochemicals that lead to unaffected plants emitting volatiles to attract predators [58], to reduce the risk of the unaffected plant being attacked while also aiding the plant being attacked. Kobayashi and Yamamura [58] specifically call this evolutionary development a form of altruism but of course there is no cognitive aspect to this altruism. These examples illustrate that evolution produces forms of sophisticated, inter-species, mixed competitive-altruistic interaction networks but without any need for a ToM, individual awareness, or strategic understanding of the interactions, and so individuals are not knowingly strategic or altruistic as we might interpret a person to be. But these evolved strategies are limited, by definition, to be fixed within the lifetime of a single agent and they can only adapt on evolutionary time-scales. This is in contrast to a ToM which allows us to adapt to multiple strategic and social contexts that may require novel solutions within the lifespan of a single person.

### v2

# An AI Theory of Mind Can Enhance Our Collective Intelligence

@@ Abstract
Collective intelligence plays a central role in many fields, from economics and evolutionary theory to neural networks and eusocial insects, and is also core to work on emergence and self-organisation in complex-systems theory. However, in human collective intelligence there is still much to understand about how specific psychological processes at the individual level give rise to self-organised structures at the social level. Psychological factors have so far played a minor role in collective-intelligence studies because the principles are often general and applicable to agents without sophisticated psychologies. We emphasise, with examples from other complex adaptive systems, the broad applicability of collective-intelligence principles, while noting that mechanisms and time scales differ markedly between cases. We review evidence that flexible collective intelligence in human social settings is improved by a particular cognitive tool: our Theory of Mind. We then hypothesise that AIs equipped with a theory of mind will enhance collective intelligence in ways similar to human contributions. To make this case, we step back from the algorithmic basis of AI psychology and consider the large-scale impact AI can have as agential actors in a “social ecology” rather than as mere technological tools. We identify several key characteristics of psychologically mediated collective intelligence and show that the development of a Theory of Mind is crucial in distinguishing human social collective intelligence from more general forms. Finally, we illustrate how individuals, human or otherwise, integrate within a collective not by being genetically or algorithmically programmed, but by growing and adapting into the socio-cognitive niche they occupy. AI can likewise inhabit one or multiple such niches, facilitated by a Theory of Mind.
@@ 1 Introduction
Collectives are capable of achieving things that individuals alone cannot. Notwithstanding the simplicity or complexity of the individuals, their aggregate behaviour can often be understood as a complex processing of information that individuals store, modify, and transfer between each other producing ‘useful’ collective behaviour at the scale of the whole collective. In most instances of Collective Intelligence (CI), where the agents might be ants in an ant colony, bees in a beehive, or neurons in a neural network, the individual is not aware of the drivers of their behaviour or the behaviour of other agents. For example, a single neuron is neither aware of its own internal processes nor that of a neuron it is connected to, nor is it aware of the end goal to which its activity contributes. Despite both this lack of awareness and the lack of a centralised controller, evolutionary and learning processes have produced an intricate, precise, and highly adaptive system that is capable of functional behaviour that would be impossible for any single neuron to achieve. In other instances of CI, such as teams of humans, or businesses interacting in economic markets, the agents themselves may be highly complex and exhibit varying degrees of purposefulness and awareness. Within this context, we draw attention to the role of psychological factors in improving the CI of human social collectives and quantifying the intelligence of social collectives, both natural and artificial. In this Introduction we review key aspects of collective intelligence and recent progress in placing AI into the collective intelligence framework.
1.1 Individual agents and their collective intelligence
@@
We can also compare the competencies of neural cells in networks to the individual competencies of ants in an ant colony. In a recent study [16] it was shown that social structures of some ant colonies are conserved between species that are separated by more than 100 million years of evolution. In the five species studied by Kay et al. [16], they found two social clusters and similarities in the division of labour that are preserved between the species. In a different study, Richardson et al. [17] showed that individuals within an ant colony play an important leadership role and that the behaviour of these individuals significantly improved the collective performance of the ants. Ants are also capable of changing their social structure in the event of pathogenic infestation of their colony. In a 2021 article, Stockmaier and colleagues [18] review the research on social distancing and other social restructuring that occurs with conspecifics in order to reduce the impact of pathogens by changing their social cues, signals, and other behaviours for the collective benefit of the colony. These two very different systems, neural networks and ant colonies, are examples of complex collective intelligences where the individuals (neurons, ants) are complex in their own right, but they signal each other in order to restructure their relationships so as to adapt their collective competencies to external signals. The neural networks are prototypical solid brains and ant colonies are prototypical liquid brains and there has been recent progress in developing collective intelligence in the context of collective adaptation [19].
@@
1.2 Recent developments using AI agents in complex adaptive systems
At the meso-scopic level, between artificial neurons and the collective behaviour of AI, is the emerging discpline of machine behaviour [36]. This field studies intelligent machines not as engineering artefacts but as a distinct class of actors with unique behavioural patterns and ecological dynamics [37]. Motivated by the increasingly pervasive integration of algorithms into human society and the challenges in formalising and predicting their complex effects, this field examines machine behaviours empirically, similar to how ethology and behavioural ecology study animals by integrating intrinsic and environmental influences [38]. In this context, it bridges micro- and macro-level questions about algorithmic interactions and their societal implications, using methods like randomised experiments and observational studies from the behavioural sciences. By drawing parallels with biological frameworks, including Tinbergen’s four dimensions of function, mechanism, development, and evolution [39], machine behaviour explores how machines interact with their environments and stakeholders, creating novel trajectories of evolution and influence that differ from organic systems. This interdisciplinary approach highlights the necessity of understanding AI’s dual role in shaping and being shaped by human systems, suggesting careful examination of its impact to harness its benefits while mitigating risks in a socio-technical context.
We can illustrate the current state of specific AI frameworks in complex adaptive systems using Large Language Models (LLMs) as an example [40]. LLMs are powerful tools for enhancing CI, but they do not yet function as autonomous agents in their own right, but instead act as mechanisms that augment human capabilities in collaboration, creativity, and decision-making. Their potential applications are highly diverse, including increasing accessibility and inclusion in online collaborations, accelerating idea generation, mediating deliberative processes, and aggregating information across groups. However, their use also introduces risks, such as disincentivising contributions to collective knowledge, fostering illusions of consensus, reducing functional diversity, and enabling the seamless production of false or misleading information. Viewed with a wide lens, LLMs not only support CI but are themselves products of it, having been trained on collective human data and refined through collective feedback. This places the use of LLMs in the context of a tool that can improve pre-existing human CI, but the LLM itself does not have agential properties in the sense that its behaviour is not produced by selecting from multiple, internally developed goals, it has very little autonomy, and it is not typically connected to a real-time working environment.
In contrast to LLMs as non-agential AIs, new work is developing cognitively informed approaches to AI-human cooperation in which they have multiple self-regarding goals that account for the goals of others, allowing them to behave with a degree of cooperative agency. In recent work by Crandall and colleagues [41], it was shown that general human–machine cooperation is achievable through a sophisticated yet fundamentally simple set of algorithmic mechanisms, with results showing that an algorithm can cooperate with humans and other machines at levels comparable to human–human cooperation in repeated stochastic games (RSGs). Through a comparison of existing algorithms, the development of a novel learning algorithm that integrates high-performing strategies with human-conducive signaling mechanisms, the algorithm (S#) was shown to form and maintain cooperative relationships across diverse RSGs. Key to its success is the inclusion of ‘cheap talk’ [42] which significantly increases mutual cooperation and highlights the importance of signaling mechanisms that are aligned with human understanding. S# also employs a variety of expert strategies that can be adapted to different partners and game types. Its expert-selection mechanism, modelled on the psychologically based recognition-primed decision-making [43], allows S# to achieve a level of flexibility and generality unmatched by traditional expert-selection methods, illustrating its potential for fostering effective cooperation in complex, real-world interactions.
Similar results have been achieved in games with more realistic complexities, approaching human levels of subtlety in social reasoning within multi-agent, mixed human-AI, environments. Work carried out on an AI playing the game Diplomacy [44] has shown that AIs can successfully defeat expert human players in complex social games of cooperation and deception without being recognised as an AI. The AI, called Cicero, integrates dialogue generation, strategic reasoning, and message filtering into its decision-making architecture in order to pass messages to other players (all of them humans who were unaware Cicero is an AI) in order to strategise about future plans in the game. Its dialogue system is based on a pre-trained language model that is further developed on human game dialogue, grounded in both the dialogue history and the game state. Dialogue generation is controlled through ‘intents’—planned actions involving Cicero and the person it is messaging—enhancing contextual relevance and strategic alignment while offloading the responsibility of learning game legality and strategy to other modules. Cicero’s strategic reasoning module employs a planning algorithm that predicts opponents’ policies based on game state and dialogue, selecting optimal actions guided by reinforcement learning with constraints to align with human-like behaviour. The approach used by Cicero has been described as an elementary form of modelling the cognitive state of other players, i.e. having a Theory of Mind. Messages are dynamically generated and filtered to ensure they are coherent, consistent with intents, and strategically sound, allowing Cicero to integrate negotiation and strategic planning in a manner indistinguishable from human players. As one of the researchers describes it [45]: “I think that the substance of what we’re doing is understanding the beliefs, goals and intentions of the other players.”. Both Cicero and S# illustrate that language combined with psychological aspects of reasoning and a strategic awareness of the environment are able to build coordination and cooperation with humans well enough for the AI to pass as a human. In order to move towards AI based more closely on human psychology [46] to enhance our collective intelligence [47] requires a greater awareness of the ecology and sociology of AI [48], and the development of interdisciplinary approaches to understanding the macroscopic consequences of these technologies.
@@ 1.2 Recent developments using AI agents in complex adaptive systems
In this article we will use information theory to quantify how much computation in a CI is ‘emergent’ and how much is simply independent information processing by single agents. In general, we wish to capture the notion of the whole (computational process) being greater than the sum of the (independent) parts. We translate this to the simple notion that to the extent to which this inequality holds: Whole - ($\sum$(∑ (Parts)>0)>0) > 0 is the extent to which we will say a system exhibits non-trivial CI, noting that there are multiple possible implementations of this approach [49]. The Parts is how much computation a single agent is carrying out from one time step to the next such that the sum is the total of all agents’ independent computations. The Whole is the totality of computation in the system, it includes all single agent computations, pairwise computations, and higher order interactions between agents. Our measure will not be unique in any of its specifics, but it serves to quantify the CI of a system for comparative analysis. This approach also has much in common with that of Moore et al. [27] in which information theory is used to measure the collective intelligence in biological systems.
@@ 1.3 The approach taken in this article
In this article we will reframe the question Watson and Levin asked in the following way: What is there in human psychology that allows us to learn to read, interpret, and re-write our interpersonal information content in a way that allows predictive control over our collective behaviour? and consider the answer in the context of pervasive, agential AI. We review some of the extensive literature showing that our Theory of Mind (ToM) is a suite of cognitive skills that allows individuals to have goal directed control over collective outcomes. Originally ToM was used to describe our ability to infer the unobserved mental states of other people [51] such as desires and beliefs, an ability humans are particularly good at and other animals much less so [52, 53]. But recently it has been shown that ToM is predictive of group performance as well [54, 55], empirically demonstrating the role of ToM in going beyond representations of the internal states of others to using that knowledge in a social setting to improve the collective outcomes for the group. In order to model ToM in a tractable fashion, we will focus on the narrower game theory of mind [56], and the Beliefs, Preferences, and Constraints (BPC) interpretation of game-theoretic decisions put forward by Gintis [57]. In this approach, what agents understand of other agents’ hidden states are the BPC that structure their observable behaviours.
@@
We will consider this question in the framework of agent interactions that extend agent utilities in a simple but novel way. We quantify our results using information theory to show the impact that a well-developed ToM has to direct agents’ behaviours in order to increase our CI. The models are simple but they illustrate the central notion that understanding the “beliefs, preferences, and constraints” [58, 59] of others can be used to improve the CI of a complex social system. The wider purpose of this work is to place recent developments, such as human-AI co-evolution [60], AI-enhanced collective intelligence [47], the collective sociology of AI and humans [48], the connection between ToM and socio-cultural niche evolution [61], and the design of intelligent cyber-physical ecosystems [62], into the context of measuring hybrid CI, CI facilitated human sociology via ToM, and how concepts from ecology, psychology, and economics can help us build a better understanding of the joint future evolution of humans and AI.
In Section 2, we describe the liquid–solid spectrum of interacting agents, review existing models of ToM, and provide perspective on the interplay between social network structures and ToM. We specifically draw attention to three key ideas that provide empirical support for either ToM or CI at the scale of human collective behaviour. First, the liquid–solid brain hypothesis establishes a cognitive spectrum from rigid and persistent to fluid and dynamical interactions between agents that exhibit collective intelligence. This approach is independent of the spatial scale of the collective, for example it can be used to study neurons, insect colonies, or human social interactions. Second, we consider the evidence for ToM, its structure at the individual agent level as a method for agents to model the unobserved control parameters that influence other agents’ behaviour. We then show how our use of language, a fluid communication channel that temporarily couples agents together in order to share information and modify behaviour, is tightly connected with our development and use of ToM. Finally, we consider the role of network topology in CI and in particular how recent work has measured the CI of social groups and related it to the individual’s capacity for ToM. In Section 3, we provide illustrative examples supporting different aspects of our argument, introducing our measure of computation and applying it to a simple empirical example. Here we note that the effectiveness of a collective depends on the macroscopic structure of interactions between agents such that any analysis will need to account for more than simple dyadic interactions and so we introduce these higher order interactions. In computing the CI between agents, a computer interacting with a monkey, we illustrate a game theory environment with an AI-biological hybrid system in which the CI value can be measured using observable behaviours. In Section 4 we review the psychology of social fluidity and the variety of social outcomes that this fluidity makes possible. We also use a simple multi-agent system to describe how a ToM can be used to improve the computational processes, i.e. the CI, of interacting agents. It is here that we connect the hidden variables (i.e. meta-parameters) argument of ToM discussed earlier with the parameters of game theory describing the beliefs, preferences and constraints that incentivise decisions [63, 64]. We also provide a formal interpretation of the higher order interactions introduced earlier in terms of game theory. Finally, in Section 6, we discuss the broader implications of this approach.
@@ 2.1 Network topology and the “Solid Brain, Liquid Brain” framework
Whole classes of models that describe which biological or artificial structures are capable of some form of computation, either in potentia or in practice, can often be usefully represented using morphospaces [78]. Computational morphospaces [7] have proven effective in studying key properties of complex adaptive systems, for example the statistical mechanics of information processing and structural variations [79, 80]. This sheds light on how energy constraints influence the evolution and adaptability of neural networks across a variety of biological systems. Arsiwalla et al. [79] have examined how liquid brains and solid brains fit within such a framework, by comparing the flexibility and adaptability of different neural architectures [79]. The dimensions of their framework are three different types of complexity that a system may display: autonomic, computational, and social complexity. We suggest that other axes for consideration are a system’s solid–liquid dimension as Solé et al. [7] and Ollé-Vila et al. [81] have done, as well as the system’s degree of ToM (see Section 2.2) and the system’s information processing capacity (see Section 3).
@@ 2.2 Models of another agent’s internal states
In the field of artificial intelligence, Jara-Ettinger [83] proposed the use of Inverse Reinforcement Learning (IRL) as a model of agential ToM, whereby agents modelling other agents’ mental states is equivalent to inferring an unobserved world model the other agent uses in their decision-making, as well as their reward function. Jara-Ettinger discusses some key limitations of IRL, such as the difficulty in recovering an agent’s beliefs and desires even while assuming that all agents are identical in their choice-making. IRL has been implemented in many different algorithmic forms, and their applicability as a basis for ToM was recently reviewed by Ruiz-Serra and Harré [84] and recent work by others in developing ToM for artificial intelligence in collaboration with people [85, 86, 87].
@@
This leads to another core finding related to ToM: the interplay between language and ToM in humans. Chomsky has noted both how well developed this ability is in humans and how communication sits in relation to our mental states [95, p. 10]:
@@
Language use is related to this causal reasoning about hidden variables via the representational view of language due to their intersection with ToM [96]. In the representational view of language people use specific grammatical structures to represent complex events and then to reason from them [97]. These structures serve as a cognitive tool, particularly in representing others’ mental states, enabling the expression of false beliefs, lies, or mistakes. It has also been shown that these are strong predictors of children’s false belief understanding, a canonical test of ToM, and targeted syntactical training can improve false belief reasoning [98]. Notably, in developmental learning, language has a stronger influence on ToM abilities than vice versa [99]. From this we identify causal reasoning about hidden cognitive states with causal reasoning about the behaviour of others via a shared representation of our social environment, mediated by specific syntactical structures.
How is language ‘causal’ though? That is, how do cognitive representations causally manifest themselves in collective outcomes via language? To some extent this might be a natural assumption to make, but its significance has been shown experimentally in that the words we choose, that grammatically represent our internal models of the world, have consequences in policymaking. In work by Thibodeau and Boroditsky [100] they have shown that two different metaphors used to describe the same numerical data regarding crime in a city: as either an infectious disease or a rampaging beast, causally influenced the likelihood of policies people chose: either preventive or punitive measures. This, they have argued [101], is the conceptual scaffolding through which we reason and that these linguistic metaphors guide our thoughts and behaviours. To date there has been little work on the levels of ToM for AI or the causal nature of language models, and this gap in the literature will need to be filled if we wish to understand how AIs can play an effective role in human collective outcomes. But if language is causal, the degree to which our language impacts outcomes is also contingent on the structural properties of the networks over which this information spreads [102]. We discuss these aspects next in the context of shared cognitive representations.
@@ 3.2 A triadic example of evolutionary learning at long time scales
This type of second order interaction occurs in other ecological examples as well [115] where Trait-Mediated Indirect Interactions (TMIIs) induce hyper-graphs of interactions between species. More complicated interactions have also been observed in which a plant that is being attacked emits infochemicals that lead to unaffected plants emitting volatiles to attract predators [114], to reduce the risk of the unaffected plant being attacked while also aiding the plant being attacked. Kobayashi and Yamamura [114] specifically call this evolutionary development a form of altruism but of course there is no cognitive aspect to this altruism. These examples illustrate that evolution produces forms of sophisticated, inter-species, mixed competitive-altruistic interaction networks but without any need for a ToM, individual awareness, or strategic understanding of the interactions, and so individuals are not knowingly strategic or altruistic as we might interpret a person to be. But these evolved strategies are limited, by definition, to be fixed within the lifetime of a single agent and they can only adapt on evolutionary time scales. This is in contrast to a ToM which allows us to adapt to multiple strategic and social contexts that may require novel solutions within the lifespan of a single person.
n the following section we illustrate how the strategic modification of an interaction network leads to improved CI, contingent on hidden cognitive variables. This provides a formal connection between hidden variables and the causal consequences of network plasticity at the level of the collective. As this is a ‘causal model’ the agent is using, it provides a useful toy model of how an AI may use cognitive tools similar to those of a human to connect the causal use of language with the consequences of network adaptation for improved collective outcomes.